Perform the indicated operation(s): 3[66 -50 -34]-2[30 -20 45]
How do I solve this equation? Thank you.

Answer by checkley77(12569)

3[66 -50 -34]-2[30 -20 45]
3*-18-2*55=-54-110=-164 ANS.